A box with a mass of 25 kg is taken to a planet where its weight is 250 N. What is the acceleration due to gravity on this planet in m/s/s?
1
Expert's answer
2020-10-23T07:14:23-0400

The weight "W" mass "m" and acceleration due to gravity "g" are connected in the following manner:


"W = mg"

Thus, expressing "g", obtain:


"g = \dfrac{W}{m}"

Substituting numbers, get:


"g = \dfrac{W}{m} = \dfrac{250N}{25kg} = 10m/s^2"

Answer. "g = 10 m\/s^2".